Suppose a bicyclist rides up a category 2 climb at 10 km/hr and rides down the same distance at a speed of 50 km/hr. Find the distance traveled if the total time on the mountain was 3 hours.
Answer by richwmiller(17219)

a+b=3,
10a=50b
a = 5/2, b = 1/2
50*.5=25
10*5/2=25
50 km total